OpenStudy (anonymous):

2x - 3y = 68 5x + y = 51 Using the two equations above, solve for y.

OpenStudy (anonymous):

5x-4,5y=170 5x+y=51 By substracting the second equasion from the first one we get -3,5y=119 5x+y=51 So y=-34 5x+y=51

OpenStudy (anonymous):

A. y = -13 B. y = -12 C. y = -14 D. y = -15

OpenStudy (trantom):

@marie123 this is parralell equations. So first you get x or y by itself for one of the equations and then you subsitute it in. 2x - 3y = 68 add 3y to both sides 2x=68 + 3y divide both sides by two x=34 + 3/2y Now subsitute all the x's in the next equation with 34 + 3/2y 5(34 + 3/2y) + y = 51 use the distributive property. 170 + 7.5y + y = 51 Now combine like terms 170 + 8.5y = 51 Subtract 170 from both sides. 8.5y = -119 Divide both sides by 8.5 y = -14 So the answer would be C. Hope this helped! :D

OpenStudy (anonymous):

thanks so much, not so good with this stuff so thank you

OpenStudy (trantom):

Just remember for problems like this you isolate one of them, then substitute in the next. :)